OpenStudy (anonymous):

Evaluate: −6 − (−4) −10 −2 24 2 I Got Two

jimthompson5910 (jim_thompson5910):

−6 − (−4) is the same as −6 + 4

OpenStudy (zzr0ck3r):

-*(-) = + -6-(-4) = -6+4 = ?

OpenStudy (anonymous):

-2

jimthompson5910 (jim_thompson5910):

And you can think of -6 + 4 as 6 - 4, but remember to change the sign of the result you get